OpenStudy (anonymous):

A triangle has vertices A(1,3,4) B (3,-1,1) C (5,1,1). What is the area of ABC

OpenStudy (anonymous):

I believe its 9.1. @amistre64 Am I correct?

OpenStudy (amistre64):

this has something to do with a dot product .... or a cross product

OpenStudy (amistre64):

assume 2d for the moment to help my brain recycle these

OpenStudy (amistre64):



OpenStudy (amistre64):

half magnitude of the cross product vector sounds familiar

OpenStudy (amistre64):

A(1,3,4) B (3,-1,1) C (5,1,1) -1-3-4 -1-3-4 -1-3-4 ----------------------------- 2,-4,-3 4,-2,-3 x 2 4 y -4 -2 z -3 -3 x = 12+6 = 18 y = -(-6+12) = -6 z = -4+16 = -12 sqrt(18^2+6^2+12^2)/2 i get about 11.22 if thats right

OpenStudy (amistre64):

6,-6,12 had an error

OpenStudy (amistre64):

sqrt(36+36+144)/2 = 7.35

OpenStudy (anonymous):

Thanks bud, but 9.1 was correct. Thanks a ton

OpenStudy (amistre64):

hmm, il have to rechk my vectors then :)

OpenStudy (amistre64):

vectors are good cross is 6, -6, 12 mag is 6sqrt(6); divided by 2 is 3sqrt(6) is 7.35 hmmm

OpenStudy (amistre64):

 \[cos(\alpha)=\frac{a.b}{|a||b|}\] \[sin(\alpha)=sin(cos^{-1}(\frac{a.b}{|a||b|}))\] \[|a||b|\frac12sin(\alpha)=|a||b|\frac12sin(cos^{-1}(\frac{a.b}{|a||b|}))\] would work too